Operations Report – April 4th

Crew 279 Operations Report 04-04-2023
SOL: 2
Name of person filing report: Ioana Dimitrova
Non-nominal systems: robotic observatory and suit number 8
Notes on non-nominal systems: Robotic observatory: awaiting for a solution from the manufacturer
Suit number 8: Problem is being assessed by crew engineer (details in Summary of suits and radios)

Spirit rover used: YES
Hours: 213.9
Beginning charge: 100%
Ending charge: 87%
Currently charging: YES
Opportunity rover used: YES
Hours: 119.7
Beginning charge: 100%
Ending charge: 89%
Currently charging: YES
Curiosity rover used: no
Hours: 226.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Perseverance rover used: no
Hours: 260.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
General notes on rovers: none
Summary of Hab operations: Music influence on stress experiment was conducted. Teamwork exercise.
WATER USE: 52.28 (see note below)
Water (static tank): 432.37
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports
Summary of suits and radios:
EVA crew reported a bit of fog forming in suit number 8 during EVA, not enough for the EVA to be dangerous or to be aborted. After testing, we noticed the suit ventilator’s airflow becomes less strong more rapidly than the others. The voltage of the battery also goes down quicker than the other two suits (suit type: helmet and suit combined type). The problem does not seem to be severe as the ventilator still works and clears the fog out but in a less optimal way than others. We suspect the battery is losing its condition. We will opt for other suits on the next EVA awaiting a repair on this one.
Radios charging nominal. Headset number 5 was replaced and labeled. It was tested and worked nominal, ready to use for the next EVA. Malfunctioning headset is in the RAM.
Summary of GreenHab operations:
Heater: Off
Supplemental light: Off
Harvest: none
Summary of ScienceDome operations: extremophiles were put in culture again because of contamination. EVA communication training. 3D printer reconstruction.
Dual split: Heat or AC, Off
Summary of RAM operations: no activity
Summary of any observatory issues: see above + cloudy weather and high winds so cupola was not opened
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support:
A smell was detected on the lower deck. We located the smell coming from above the airlock that leads to the other parts of the station. We checked to see if the smell was coming from gas with the gas detector but the gas detector did not detect anything. We then tried touching and observing to see if there were any burnt cables or plastic caused eventually by a short circuit. After that, we called Mission Support who responded that this smell could occur, coming from the exhaust of the water heater when there are high winds.

Copyright © The Mars Society. All rights reserved. | Main Site